factoring trinomials
1
Expert's answer
2014-10-16T12:37:17-0400
This problem is related to finding the real roots of equation ax2+bx+c=0. If b2-4ac>0, then equation has two distinct roots x1, x2 and factorization is expressed by ax2+bx+c=a(x-x1)(x- x2). If b2-4ac=0, then equation has two equal roots
x1= x2 =z and factorization is expressed by ax2+bx+c=a(x-z)2. If b2-4ac<0, then equation has no real roots and it is impossible to factor trinomial ax2+bx+c.